OpenStudy (dj3strella):

Help 4. What is the completed factored form of 5x^2 – 45? 5(x + 3)(x – 3) 5(x – 3)^2 5(x^2 – 9) 5(x + 9)(x – 9)

OpenStudy (anonymous):

Tell us what you think so we can see where you got stuck

OpenStudy (dj3strella):

I dont know how to solve the problem @Crixpack

OpenStudy (anonymous):

okay, but do you understand factorizing in general?

OpenStudy (dj3strella):

yes

OpenStudy (anonymous):

okay, can you spot any common factors between 5x^2 and 45?

OpenStudy (helder_edwin):

\[\large 5x^2-45=\color{red}{5}x^2-\color{red}{5}\cdot9=5(x^2-9)= 5(x+3)(x-3). \]